OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –5? A. y = –3x – 5 B. y = 3x + 5 C. y = –3x + 5 D. y = 3x – 5

OpenStudy (joannablackwelder):

y = mx + b m is slope b is y-intercept

OpenStudy (anonymous):

is it a or b

OpenStudy (joannablackwelder):

y = (-3)x + (-5) And, yes, it does simplify to A

OpenStudy (anonymous):

Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (y – 1) = 3(x + 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (joannablackwelder):

Point-slope form is y - y1 = m(x - x1)

OpenStudy (joannablackwelder):

Where slope is m and a point it passes through is (x1,y1)

OpenStudy (joannablackwelder):

Any ideas, @drewsean ?